# In constructing our theory of cosmology, our inspiration is the double-cone geometry of [2]. In the context of two-dimensional (Jackiw–Teitelboim) gravity, the metric of the double cone is that of AdS2 metric in Rindler (i.e., black hole) coordinates:

### v1

@@ 1 Introduction
In constructing our theory of cosmology, our inspiration is the double-cone geometry of [2]. In the context of two-dimensional (Jackiw–Teitelboim) gravity, the metric of the double cone is that of AdS2 metric in Rindler (i.e., black hole) coordinates:
@@
For this to work, one needs to better understand the singularity at $u=0$. One approach is to resolve the singularity by complexifying the metric slightly; we will review this complexification (and others) in the next section. When one does so, the two asymptotic regions are connected by a smooth (although complex) geometry near $u=0$. The important point is that there is a clear criterion, known as the Kontsevich–Segal (KS) condition, that characterizes the precise circumstances under which a complex geometry can be viewed as a perturbatively stable saddle point [4, 5]. This complexification procedure allows one to understand exactly which perturbative fluctuations are allowed in the double cone geometry [6]. [^3] Indeed, the role of complex saddle points has been recognized elsewhere, including rotating black holes and cosmology [7, 8, 9], global symmetries in two dimensions [10], JT gravity in nearly dS [11], and holographic Schwinger–Keldysh contours [12, 13, 14]. Most recently, this includes constraints on inflation [15, 16] and more general aspects of cosmology [17, 18].
@@
There are a number of antecedents for this construction: the Hartle–Hawking no-boundary proposal constructs cosmological saddle points by connecting a smooth compact Euclidean section to a Lorentzian cosmology [7]. A similar recent construction with a negative cosmological constant connects a Euclidean AdS space with a boundary to a Lorentzian cosmology [22]. Our construction is different in that these constructions connect a Euclidean solution to a Lorentzian one at a moment of time symmetry in the cosmology, while our construction will connect two Lorentzian solutions close to the singularity in the cosmology. As in the previous constructions, these saddle points could be used to define a wavefunction on a spatial slice of the cosmology. Our focus, however, is on understanding their description from the dual CFT perspective.
This CFT interpretation is possible because the solutions have two Lorentzian asymptotically AdS boundaries, as in the double cone. So, they can be interpreted in terms of the statistics of the density of states in the dual CFT. We will consider this in AdS2 and AdS3. In the AdS2 case, our cosmologies are interpreted as additional contributions to the usual spectral form factor. In AdS3, however, they have a more interesting interpretation. They give contributions to a novel observable that describes correlations between the spectrum at different values of the energy. Our main result is that the torus cosmology contributes to the spectral observable
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta E}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu,$ (1.4)
@@
In the next section, we will describe in detail our geometries, and their complexification, and analyze the Kontsevich–Segal condition which guarantees their stability. This will lead us to the cosmological interpretation of our new spectral observable (1.4). In section 3 we will discuss the action and one-loop determinants on these solutions and make a specific prediction for this observable. In section 4 we analyze field theory in this background, which is described in terms of the spectrum of black hole (anti-)quasinormal modes. We conclude in section 5 with some speculation and discussion of future directions. Appendix A contains comments on the higher-dimensional cases.
@@ 2 Cosmology from complex geometries
This coordinate is sometimes introduced as a convenient way to describe the whole Einstein–Rosen bridge, as $u\in(-\infty,\infty)$ covers the two exterior regions. It thus provides a good coordinate system for the double-cone geometry. In the cosmology, since $f(r)<0$, $u$ becomes imaginary, and the cosmology inside the horizon corresponds to (a portion of) the imaginary axis in the complex $u$-plane.
@@
Of course, not every complex metric will be allowed. We will demand that our metrics obey the Kontsevich–Segal criterion [4, 5], which requires that the real part of the matrix $\sqrt{\det g}g^{\mu\nu}$ be positive definite. For a diagonal metric $g_{\mu\nu}=\lambda_{\mu}\delta_{\mu\nu}$, this is the same as the statement that the sum of the complex phases of the metric components is less than $\pi$:
@@
Metrics that satisfy the Kontsevich–Segal criterion have the property that perturbative field theories based on form fields are well defined, in the sense that all of the Gaussian integrals that appear when one expands a field theory around this background are convergent. So it appears to be a necessary (though possibly not sufficient) criterion for a metric to be treated as a well-defined and perturbatively stable saddle-point contribution to a gravitational path integral.
This will constrain significantly the possible choices of contour, in a way that depends non-trivially on dimension. In the remainder of this section, we will study the allowed contours in two and three dimensions; we relegate a discussion of higher dimensions to an appendix. In section 2.1, we study two dimensions, where the contours include a Hartle–Hawking-like contour which interpolates between a Euclidean and a Lorentzian geometry, and new contours that contribute to the spectral form factor, interpolating between Lorentzian geometries. In section 2.2, we study three dimensions, where we find contours that satisfy the Kontsevich–Segal criterion and interpolate between two Lorentzian cones whose cycles are swapped compared to each other. These are the main results of our paper.
@@ 2.1 Warm-up: new contours in two dimensions
The double-cone geometry of [2, 6] is obtained by identifying $t\sim t+T$ and considering real $u\in(-\infty,\infty)$. Since this geometry is real and Lorentzian, it saturates the Kontsevich–Segal bound. We can then regulate the geometry by taking $u=\rho+\mathrm{i}\epsilon$. The resulting geometry then obeys the Kontsevich–Segal condition [5].
In the complex $u$-plane, we have other interesting solutions: if we set $u=\rho\pm\mathrm{i}\frac{\pi}{2}$, we find a Euclidean AdS2 metric. As a real Euclidean metric, this of course also obeys the Kontsevich–Segal condition. This was exploited in [6] to understand the modified time-translation generator associated with the version of the double cone, where we go around the singularity at $\rho=0$ in the complex plane. Similarly, if we set $u=\rho\pm\mathrm{i}\pi$, we obtain the Lorentzian AdS2 space-time again.
Our primary interest, however, is the solution where $u$ is purely imaginary: $u=\mathrm{i}\rho$. This describes a cosmology that runs from a big bang at $\rho=0$, reaching a maximum size where it crosses the Euclidean section at $\rho=\pm\frac{\pi}{2}$, and ending with a big crunch at $\rho=\pm\pi$. As this geometry saturates the Kontsevich–Segal criterion, we would like to understand whether it can be regulated in the same way as the double-cone geometry.
@@
Let us consider a general contour $u(\rho)$, and write $u=x+\mathrm{i}y$. The Kontsevich–Segal criterion (2.5) becomes
@@
These are plotted for $y\in(0,\frac{\pi}{2})$ in figure 2. The contours are symmetric about $y=\frac{\pi}{2}$ so the contours for $y\in(\frac{\pi}{2},\pi)$ are obtained by reflection. As expected, we see that the direction of the contour is unrestricted around the Euclidean AdS2 at $y=\frac{\pi}{2}$. The critical contours that start at the origin return to $(x,y)=(0,\pi)$, while the ones that come in from infinity go back out to infinity along $y=0$ and $y=\pi$. We see that the latter have $y\to 0$ as $x\to\infty$ for all values of $a$.
@@
Figure 2: Critical contours which saturate the Kontsevich–Segal condition, plotted in the complex $u$-plane for the AdS2 case.
A familiar example of a contour involving the cosmology is a Hartle–Hawking-like contour shown in figure 1(b), where we run along the Euclidean solution at $y=\pm\frac{\pi}{2}$, and then along a portion of the imaginary axis, defining a state in the cosmology.
We are interested instead in contours that asymptotically approach the cone geometries. We note that there is a Lorentzian cone geometry at $y=0$, but there is also another one at $y=\pm\pi$, and we can consider a contour that runs between them. One option for such a contour is the one pictured in figure 1(a), running through the cosmology. There is, however, a family of other possibilities consistent with the Kontsevich–Segal criterion.
From the boundary perspective, these contours have the same asymptotics as the double cone, so they correspond to additional contributions to |$\mkern 1.3mu\overline{\mkern-1.3mu\absolutevalue{Y_{E,\Delta E}(T)}^{2}\mkern-1.3mu}\mkern 1.3mu$. [^5] We can also consider contours which start from $y=0$ at large $x$, turn vertical at $y=\mathrm{i}\frac{\pi}{2}$, and return to $x\to\infty$ as $y\to\pi$. These again have the same asymptotics and therefore represent additional contributions to |$\mkern 1.3mu\overline{\mkern-1.3mu\absolutevalue{Y_{E,\Delta E}(T)}^{2}\mkern-1.3mu}\mkern 1.3mu$. It is somewhat surprising that we have additional contributions to this observable, since the result from the double cone matched nicely with the expectations for this from the dual theory. It would be interesting to better understand what the role of these additional contributions is and to see if we could disentangle some distinctive signatures related to the cosmology. Unfortunately, our understanding of how to evaluate the gravitational path integral when multiple saddles contribute is limited. In the next section, we will see that in three dimensions there is a saddle point involving the cosmology, which contributes to a distinct observable, which should make probing the role of the cosmology more straightforward in that context.
@@ 2.2 Contours in three dimensions
with $\tilde{t}\sim\tilde{t}+\tilde{T}$ and $\tilde{\phi}\sim\tilde{\phi}+2\pi$, where $\tilde{T}=\frac{4\pi^{2}}{T}$. This metric saturates the Kontsevich–Segal criterion but can be regulated as before by subtracting a small imaginary part from $u$. We conclude that the contour $u=\tilde{\rho}+\mathrm{i}\frac{\pi}{2}$ contributes to
|$\mkern 1.3mu\overline{\mkern-1.3mu\absolutevalue{Y_{\tilde{E},\Delta E}\quantity(\frac{4\pi^{2}}{T})}^{2}\mkern-1.3mu}\mkern 1.3mu,\qquad\text{with}\qquad\tilde{E}=\frac{4\pi^{2}r_{\mathrm{h}}^{2}}{8GT^{2}}.$ (2.15)
@@
Of course, this geometry exactly saturates the Kontsevich–Segal condition.
Our goal is to deform this cosmological contour to satisfy the Kontsevich–Segal criterion. We construct a contour that approaches the horizontal contours asymptotically and passes through the cosmology, as sketched in figure 3. This contour will offer some insight into the microscopic description of the cosmology.
@@
To investigate the Kontsevich–Segal criterion, we again write $u=x+\mathrm{i}y$ with $y\in\quantity(0,\frac{\pi}{2})$, and the condition (2.5) becomes
@@
plotted in figure 4. These critical contours are horizontal at large $x$, with the curves with $\frac{\mathrm{d}y}{\mathrm{d}x}>0$ approaching the origin as $x\to 0$, and the curves with $\frac{\mathrm{d}y}{\mathrm{d}x}<0$ approaching $(x,y)=\quantity(0,\frac{\pi}{2})$.
@@
Figure 4: Limiting contours in the complex $u$-plane in three dimensions.
@@
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta E}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu,$
@@ 3 Gravity on complexified geometries
We will focus in this section on the more interesting case of three dimensions, where the action is the usual Einstein–Hilbert action, but a similar analysis would apply to the two-dimensional case, using the JT gravity action.
3.1 Classical Einstein–Hilbert action
@@
as the integrand is an odd function of $\rho$. Similarly defining the volume measure in the boundary term to be $\sqrt{-h}=r_{\mathrm{h}}^{2}\sinh\rho\cosh\rho$, we see that the two boundary terms will cancel, so that the overall value of the action vanishes. [^7] It might not seem so clear that such an analytic continuation prescription is natural for the boundary terms, as these are only evaluated at the endpoints, not along the contour. However, the prescription for defining the boundary terms must be chosen consistently with the bulk prescription for the action to remain stationary under variations preserving the relevant boundary conditions. This agrees with the claim in [2] that the action should in general vanish for the double cone, as the two sides are complex conjugates of each other. Note that this is different from simply taking the action (3.4) of the cone region outside the horizon twice, which is what we might be naturally led to do if we ignored the deformation of the contour near $\rho=0$.
@@
In the boundary terms, we have $\sqrt{h}=\frac{1}{2}\sinh 2u$; at $u=-\rho_{\mathrm{max}}+\mathrm{i}\frac{\pi}{2}$ we get $\sinh 2u=\sinh 2\rho_{\mathrm{max}}$, so the boundary terms at each end have the same sign in this contour, and the boundary contributions cancel the divergent part of the bulk, leaving us with zero action.

### v3

# In constructing our theory of cosmology, our inspiration is the double-cone geometry of [2]. In the context of two-dimensional (Jackiw-Teitelboim) gravity, the metric of the double cone is that of AdS2 metric in Rindler (i.e., black hole) coordinates:

@@ 1 Introduction
In constructing our theory of cosmology, our inspiration is the double-cone geometry of [2]. In the context of two-dimensional (Jackiw-Teitelboim) gravity, the metric of the double cone is that of AdS2 metric in Rindler (i.e., black hole) coordinates:
@@
For this to work, one needs to better understand the singularity at $u=0$. One approach is to resolve the singularity by complexifying the metric slightly; we will review this complexification (and others) in the next section. When one does so, the two asymptotic regions are connected by a smooth (although complex) geometry near $u=0$. The important point is that there is a clear criterion, known as the Kontsevich-Segal (KS) condition, that characterizes the precise circumstances under which a complex geometry can be viewed as a perturbatively stable saddle point [4, 5]. This complexification procedure allows one to understand exactly which perturbative fluctuations are allowed in the double cone geometry [6]. [^3] Indeed, the role of complex saddle points has been recognized elsewhere, including rotating black holes and cosmology [7, 8, 9], global symmetries in two dimensions [10], JT gravity in nearly dS [11], and holographic Schwinger-Keldysh contours [12, 13, 14]. Most recently, this includes constraints on inflation [15, 16] and more general aspects of cosmology [17, 18].
@@
There are a number of antecedents for this construction: the Hartle-Hawking no-boundary proposal constructs cosmological saddle points by connecting a smooth compact Euclidean section to a Lorentzian cosmology [7]. A similar recent construction with a negative cosmological constant connects a Euclidean AdS space with a boundary to a Lorentzian cosmology [22]. Our construction is different in that these constructions connect a Euclidean solution to a Lorentzian one at a moment of time symmetry in the cosmology, while our construction will connect two Lorentzian solutions close to the singularity in the cosmology. As in the previous constructions, these saddle points could be used to define a wavefunction on a spatial slice of the cosmology. Our focus, however, is on understanding their description from the dual CFT perspective.
This CFT interpretation is possible because the solutions have two Lorentzian asymptotically AdS boundaries, as in the double cone. So, they can be interpreted in terms of the statistics of the density of states in the dual CFT. We will consider this in AdS2 and AdS3. In the AdS2 case, we can construct contours that satisfy the KS condition, but they will not satisfy the same boundary conditions for the dilaton at both ends, so they are not contributions to a spectral form factor in a single theory. [^5] We thank Martin Sasieta for pointing out the importance of the dilaton here. In AdS3, the new contours give contributions to a novel observable that describes correlations between the spectrum at different values of the energy in a single theory. Our main result is that the torus cosmology contributes to the spectral observable
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta\tilde{E}}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu,$ (1.4)
@@
In the next section, we will describe in detail our geometries, and their complexification, and analyze the Kontsevich-Segal condition which guarantees their stability. This will lead us to the cosmological interpretation of our new spectral observable (1.4). In section 3 we will discuss the action and one-loop determinants on these solutions and make a specific prediction for this observable. In section 4 we analyze field theory in this background, which is described in terms of the spectrum of black hole (anti-)quasinormal modes. We conclude in section 5 with some speculation and discussion of future directions. Appendix A contains comments on the higher-dimensional cases.
@@ 2 Cosmology from complex geometries
This coordinate is sometimes introduced as a convenient way to describe the whole Einstein-Rosen bridge, as $u\in(-\infty,\infty)$ covers the two exterior regions. It thus provides a good coordinate system for the double-cone geometry. In the cosmology, since $f(r)<0$, $u$ becomes imaginary, and the cosmology inside the horizon corresponds to (a portion of) the imaginary axis in the complex $u$-plane.
@@
Of course, not every complex metric will be allowed. We will demand that our metrics obey the Kontsevich-Segal criterion [4, 5], which requires that the real part of the matrix $\sqrt{\det g}g^{\mu\nu}$ be positive definite. For a diagonal metric $g_{\mu\nu}=\lambda_{\mu}\delta_{\mu\nu}$, this is the same as the statement that the sum of the complex phases of the metric components is less than $\pi$:
@@
Metrics that satisfy the Kontsevich-Segal criterion have the property that perturbative field theories based on form fields are well defined, in the sense that all of the Gaussian integrals that appear when one expands a field theory around this background are convergent. So it appears to be a necessary (though possibly not sufficient) criterion for a metric to be treated as a well-defined and perturbatively stable saddle-point contribution to a gravitational path integral.
This will constrain significantly the possible choices of contour, in a way that depends non-trivially on dimension. In the remainder of this section, we will study the allowed contours in two and three dimensions; we relegate a discussion of higher dimensions to an appendix. In section 2.1, we study two dimensions, where the contours include a Hartle-Hawking-like contour which interpolates between a Euclidean and a Lorentzian geometry, and new contours that contribute to the spectral form factor, interpolating between Lorentzian geometries. In section 2.2, we study three dimensions, where we find contours that satisfy the Kontsevich-Segal criterion and interpolate between two Lorentzian cones whose cycles are swapped compared to each other. These are the main results of our paper.
@@ 2.1 Warm-up: new contours in two dimensions
The double-cone geometry of [2, 6] is obtained by identifying $t\sim t+T$ and considering real $u\in(-\infty,\infty)$. Since this geometry is real and Lorentzian, it saturates the Kontsevich-Segal bound. We can then regulate the geometry by taking $u=\rho+\mathrm{i}\epsilon$. The resulting geometry then obeys the Kontsevich-Segal condition [5].
In the complex $u$-plane, we have other interesting solutions: if we set $u=\rho\pm\mathrm{i}\frac{\pi}{2}$, we find a Euclidean AdS2 metric. As a real Euclidean metric, this of course also obeys the Kontsevich-Segal condition. This was exploited in [6] to understand the modified time-translation generator associated with the version of the double cone, where we go around the singularity at $\rho=0$ in the complex plane. Similarly, if we set $u=\rho\pm\mathrm{i}\pi$, we obtain the Lorentzian AdS2 space-time again.
Our primary interest, however, is the solution where $u$ is purely imaginary: $u=\mathrm{i}\rho$. This describes a cosmology that runs from a big bang at $\rho=0$, reaching a maximum size where it crosses the Euclidean section at $\rho=\pm\frac{\pi}{2}$, and ending with a big crunch at $\rho=\pm\pi$. As this geometry saturates the Kontsevich-Segal criterion, we would like to understand whether it can be regulated in the same way as the double-cone geometry.
@@
Let us consider a general contour $u(\rho)$, and write $u=x+\mathrm{i}y$. The Kontsevich-Segal criterion (2.5) becomes
@@
These are plotted for $y\in(0,\frac{\pi}{2})$ in figure 2. The contours are symmetric about $y=\frac{\pi}{2}$ so the contours for $y\in(\frac{\pi}{2},\pi)$ are obtained by reflection. As expected, we see that the direction of the contour is unrestricted around the Euclidean AdS2 at $y=\frac{\pi}{2}$. The critical contours that start at the origin return to $(x,y)=(0,\pi)$, while the ones that come in from infinity go back out to infinity along $y=0$ and $y=\pi$. We see that the latter have $y\to 0$ as $x\to\infty$ for all values of $a$. The critical contours correspond to the case where the KS criterion is saturated. Hence, a physical contour’s slope at each point should be bounded above and below by two critical contours to satisfy the KS criterion.
@@
Figure 2: Critical contours which saturate the Kontsevich-Segal condition, plotted in the complex $u$-plane for the AdS2 case.
A familiar example of a contour involving the cosmology is a Hartle-Hawking-like contour shown in figure 1(b), where we run along the Euclidean solution at $y=\pm\frac{\pi}{2}$, and then along a portion of the imaginary axis, defining a state in the cosmology.
We are interested instead in contours that asymptotically approach the cone geometries. We note that there is a Lorentzian cone geometry at $y=0$, but there is also another one at $y=\pm\pi$, and we can consider a contour that runs between them. One option for such a contour is the one pictured in figure 1(a), running through the cosmology. There is, however, a family of other possibilities consistent with the Kontsevich-Segal criterion.
From the boundary perspective, these contours have the same metric asymptotics as the double cone. However, if we consider these as solutions in JT gravity, they have different asymptotics for the dilaton; (2.6) is a solution with $\phi=\phi_{\mathrm{h}}\cosh u$, so if the dilaton has a positive value at real $u$ it will have a negative value for $u=\rho+\mathrm{i}\pi$. Thus, these contours do not contribute to the spectral form factor in JT gravity. [^6] We thank Martin Sasieta for pointing out this issue with the dilaton.
@@ 2.2 Contours in three dimensions
with $\tilde{t}\sim\tilde{t}+\tilde{T}$ and $\tilde{\phi}\sim\tilde{\phi}+2\pi$, where $\tilde{T}=\frac{4\pi^{2}}{T}$. This metric saturates the Kontsevich-Segal criterion but can be regulated as before by subtracting a small imaginary part from $u$. We conclude that the contour $u=\tilde{\rho}+\mathrm{i}\frac{\pi}{2}$ contributes to
|$\mkern 1.3mu\overline{\mkern-1.3mu\absolutevalue{Y_{\tilde{E},\Delta E}\quantity(\frac{4\pi^{2}}{T})}^{2}\mkern-1.3mu}\mkern 1.3mu,\qquad\text{with}\qquad\tilde{E}=\frac{T^{2}r_{\mathrm{h}}^{2}}{32\pi^{2}G}.$ (2.15)
@@
Of course, this geometry exactly saturates the Kontsevich-Segal condition.
Our goal is to deform this cosmological contour to satisfy the Kontsevich-Segal criterion. We construct a contour that approaches the horizontal contours asymptotically and passes through the cosmology, as sketched in figure 3. This contour will offer some insight into the microscopic description of the cosmology.
@@
To investigate the Kontsevich-Segal criterion, we again write $u=x+\mathrm{i}y$ with $y\in\quantity(0,\frac{\pi}{2})$, and the condition (2.5) becomes
@@
plotted in figure 4. These critical contours are horizontal at large $x$, with the curves with $\frac{\mathrm{d}y}{\mathrm{d}x}>0$ approaching the origin as $x\to 0$, and the curves with $\frac{\mathrm{d}y}{\mathrm{d}x}<0$ approaching $(x,y)=\quantity(0,\frac{\pi}{2})$. Similar to the previous case, the allowed contours lie in between the critical contours, with their slope at each point bounded above and below by the critical contours.
@@
Figure 4: Critical contours in the complex $u$-plane in three dimensions.
@@
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta\tilde{E}}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu,$
@@ 3 Gravity on complexified geometries
We will focus in this section on the more interesting case of three dimensions, where the action is the usual Einstein-Hilbert action, but a similar analysis would apply to the two-dimensional case, using the JT gravity action.
3.1 Classical Einstein-Hilbert action
@@
as the integrand is an odd function of $\rho$. Similarly, defining the volume measure and extrinsic curvature in the boundary term to be $\sqrt{-h}=r_{\mathrm{h}}^{2}\sinh\rho\cosh\rho$ and $\sqrt{-h}K=r_{\mathrm{h}}^{2}\cosh 2\rho$ respectively, we see that two boundary terms will also cancel so that the overall value of the action vanishes. [^8] It might not seem so clear that such an analytic continuation prescription is natural for the boundary terms, as these are only evaluated at the endpoints, not along the contour. However, the prescription for defining the boundary terms must be chosen consistently with the bulk prescription for the action to remain stationary under variations preserving the relevant boundary conditions. This agrees with the claim in [2] that the action should in general vanish for the double cone, as the two sides are complex conjugates of each other. Note that this is different from simply taking the action (3.4) of the cone region outside the horizon twice, which is what we might be naturally led to do if we ignored the deformation of the contour near $\rho=0$.
@@
In the boundary terms, we have $\sqrt{h}=\frac{1}{2}\sinh 2u$; at $u=-\rho_{\mathrm{max}}+\mathrm{i}\frac{\pi}{2}$ we get $\sinh 2u=\sinh 2\rho_{\mathrm{max}}$, so the boundary terms at each end have the same sign in this contour, and the boundary contributions cancel the divergent part of the bulk, leaving us with zero action. In fact, this is in line with the intuition that the classical action would contribute as an overall phase of the form $\mathrm{e}^{\mathrm{i}\tilde{E}\tilde{T}-\mathrm{i}ET}$ to the observable (2.21): since $\tilde{E}\tilde{T}=ET$ in this geometry, the phase is zero.